My math problem states:
{9x + 9y=0}
{5x - 2y= 14}
The instructions are to solve the system of equations by graphing, but I am stuck on how to solve the problem. I have tried several ways but neither equal correctly, please help :-)

Answer by mananth(16946)

9 x + 9 y = 0 .............1

5 x -2 y = 14 .............2
Find the x & y intercept points by plugging x=0,y=0 in both equations.
equation I
when x= 0 y= 0 ,( 0 , 0 )
when y = 0 x= 0 ,( 0 , 0 )
x=1,y=-1,(1,-1)
Equation 2
when x= 0 y= -7 ( 0 , -7 )
when y = 0 x= 2.8 ( 0 , 2.8 )


(2,-2) is the solution
m.ananth@hotmail.ca